# Mitigating Multimodal Hallucinations via Gradient-based Self-Reflection

### v2

@@ Appendix F Results on MM-Vet
Appendix G Results on ScienceQA and Vizwiz
We evaluate our method on two complementary multimodal benchmarks. ScienceQA Lu et al. (2022) integrates images, textual context, and curriculum knowledge, requiring models to perform structured multimodal reasoning. VizWiz Gurari et al. (2018), in contrast, consists of visual questions collected from blind users and features real-world challenges such as low-quality images, conversational queries, and unanswerable cases. These datasets jointly assess both reasoning under structured multimodal contexts and robustness in unconstrained real-world settings. As shown in Table 10, our approach consistently improves over the LLaVA-1.5 baseline. These gains demonstrate the effectiveness of our hallucination mitigation strategy in enhancing visual grounding across both knowledge-driven and real-world VQA tasks.
@@ Appendix G Results on ScienceQA and Vizwiz
Appendix H Revisiting the Accuracy–Informativeness Trade-off
MLLM Method $C_{s}$ $\downarrow$ $C_{i}$ $\downarrow$ R $\uparrow$ Len $\uparrow$
LLaVA-v1.5 base 48.8 13.4 78.6 99.8
PAI Liu et al. (2024c) 24.8 6.9 - -
Middle Jiang et al. (2025) 25.0 6.7 - -
Ours_ES_12% 23.5 6.5 55.0 54.1
Appendix I Results on MME
Our evaluation on MME Fu et al. (2024) dataset is presented in Tab. 12. Our method achieves better overall (Total) results and equal or improved performance in existence and counting, demonstrating its effectiveness in object recognition. However, it does not improve position accuracy and exhibits varying behavior on color. This diversity may stem from the inherent capabilities of MLLMs, which cannot be solely enhanced through token influence balancing.
@@ Appendix I Results on MME
Appendix J Other Results of POPE
Table 13: More Results on POPE Li et al. (2023b).
@@
Appendix K Different Sampling Strategies
Tab. 14 presents an ablation study on sampling strategies (non-greedy vs. greedy). We follow the non-greedy sampling setting of VCD Leng et al. (2024), where both top-p and temperature are set to 1. As shown, our method consistently improves performance across both sampling strategies.
@@
Appendix L Different Model Size
@@
Appendix M Gradient Computation Details and Efficiency Ablation
Table 16: Ablation Study on Gradient Methods on the POPE MSCOCO Adversarial Dataset
Methods Accuracy F1 Average Speed(ms)
MPLUG-Owl2 IG Enguehard (2023) 83.4 82.9 20335
direct 84.2 83.7 385
Our method obtains gradients directly through PyTorch’s ‘torch.autograd.grad’ on input tokens, eliminating the need for manual derivations and facilitating straightforward reproduction. For comparison, we evaluate Integrated Gradients (IG) Enguehard (2023); Lundstrom et al. (2022); Kapishnikov et al. (2021) using the SIG Enguehard (2023) implementation; Table 16 presents this ablation. In the table, “IG” denotes the SIG-based results, while “direct” refers to our torch.autograd.grad approach. Both methods yield comparable accuracy and F1‐score, but the direct‐gradient variant is substantially more efficient.
Appendix N Norm Selection for Token Influence
[…]
Table 17: Norm Strategies on POPE MSCOCO Adversarial Dataset.
@@
L1 83.5 82.1 82.5 82.1 84.2 83.7
L2 83.2 81.9 79.5 79.6 83.2 82.9
L 83.4 82.0 82.1 81.8 80.8 80.6
Appendix O Hyper parameter Study
Maximum $\alpha_{m}$. To determine the optimal maximum amplification factor ($\alpha_{m}$) and understand its impact on the model performance, we conducted a search over values from 1 to 6 using the LLaVA-v1.5 model on the POPE dataset for discriminative tasks. For open-ended generation tasks on a subset of the MSCOCO dataset following Deng et al. (2024), we observed garbled text when $\alpha_{m}$ was set to 5; therefore, we limited our search to values from 1 to 4. As shown in Tab. 18 and Tab. 19, discriminative task on POPE is less sensitive to the value of $\alpha_{m}$. The performance on the generative task gets improved while $\alpha_{m}$ increases but later drops. Therefore, in our experiments, the optimal maximum amplification factor ($\alpha_{m}$) is set to 5 and 3 for discriminative and generation tasks, respectively.
Table 18: $\alpha_{m}$ Study For Discriminative Task On POPE Li et al. (2023b) in MSCOCO Adversarial Setting.
@@
Early Stopping Threshold. To set the early stopping threshold properly, we conducted a search on a subset of the MSCOCO dataset subset following Deng et al. (2024). Recognizing that the visual influence ratio varies across models, we first analyze the sample-wise visual influence to identify an appropriate range for this study. By searching over the corresponding range, we show results in Tab. 20. These results demonstrate that varying the ES threshold primarily mediates the trade-off between recall and hallucination rate. Our goal is to have balanced recall ($R$) and instance-level hallucination ($C_{I}$), leading us to select thresholds of $7\%$ for LLaVA-v1.5 and LLaVA-v1.6, $25\%$ for InstructBLIP, $2.5\%$ for mPLUG-Owl2 and $10\%$ for InternVL2. We additionally ran an experiment to measure the ES activation rate using LLaVA-v1.5 with ES threshold $7\%$. As shown in Tab. 21, ES fires on only $8.7\%$ of the test samples, and when it does, the generated responses are on average just 0.7 tokens shorter. This indicates that ES rarely, and only minimally, truncates outputs.
@@
Appendix P Confidence and Visual Influence
@@
Appendix Q Question Category Results on the AMBER Dataset
@@
Appendix R Computational Cost
In this section, we analyze the computational cost of the proposed method. On the POPE dataset, our method increases the average computational cost by $101.44\%$ in TFLOPs. This is expected for decoding-based approaches, as they require two forward passes. Importantly, the visual encoder is executed only once, and the second pass processes a limited number of input tokens. Thus, although additional computation is introduced, the overall overhead remains within a practical range.
Method TFLOPs Relative Increase
LLaVA-v1.5 $9.68$
Ours $19.49$ $101.44\%$
Appendix S Influence Ratio in VQA
@@
Appendix T Influence Ratio in Image Caption
@@
Appendix U Qualitative Example on Occluded Images
We include a qualitative example in Fig. 9, where a sedan and a building are partially occluded by a white truck, our method prevents the baseline model from hallucinating of persons and vehicles behind the occluding object. This demonstrates our method remains effective on images consisting of occlusions.
@@
Image caption w/o Ours: The image features a white truck parked on a city street, with graffiti covering its side. The truck is positioned near a crosswalk, and there are several other vehicles in the scene, including a car and a bus. In addition to the vehicles, there are a few people walking around the area. Image caption w Ours: The image features a large white truck parked on a city street, with graffiti covering its side. The truck is positioned near a crosswalk, and there is another vehicle visible in the background. Additionally, there are buildings in the scene, suggesting an urban setting.
Appendix V Additional Implementation and Experimental Details
Noun tokens are identified using the spaCy library via its en_core_web_sm model. For experiments on the Amber dataset Wang et al. (2023), we adopt the original data splits and evaluation metrics. In the MSCOCO Lin et al. (2014) subset, we follow the data partitioning and evaluation protocol of Deng et al. Deng et al. (2024), with splits available in their official repository. For the LLaVA-QA90 Liu et al. (2024b), MME Fu et al. (2024) and POPE Li et al. (2023b) datasets, our setup replicates that of Leng et al. Leng et al. (2024) and use their provided scoring scripts for LLaVA-QA90. Experiments on MMBench Liu et al. (2025), MM-Vet Yu et al. (2023) follow the VLMEvalKit_InternVL2_5 repository. All comparison methods are executed using their official code; we only modify them to enforce greedy sampling and a uniform maximum generation length to align with our experimental settings.
Appendix W Limitations and Future Work
Our method is limited to white-box MLLMs, as it requires access to gradients. Its effectiveness depends on the baseline MLLM’s original visual influence ratio, and the importance of visual information. As a post-processing technique, our method does not involve model training. In future work, we aim to explore how insights from GACD can guide and improve training strategies for enhanced visual perception in MLLMs.
Appendix X Broader Impacts

### v3

@@ 6 Results on MM-Vet
7 Results on ScienceQA and Vizwiz
We evaluate our method on two complementary multimodal benchmarks. ScienceQA [lu2022learn] integrates images, textual context, and curriculum knowledge, requiring models to perform structured multimodal reasoning. VizWiz [gurari2018vizwiz], in contrast, consists of visual questions collected from blind users and features real-world challenges such as low-quality images, conversational queries, and unanswerable cases. These datasets jointly assess both reasoning under structured multimodal contexts and robustness in unconstrained real-world settings. As shown in Table 5, our approach consistently improves over the LLaVA-1.5 baseline. These gains demonstrate the effectiveness of our hallucination mitigation strategy in enhancing visual grounding across both knowledge-driven and real-world VQA tasks.
@@ 7 Results on ScienceQA and Vizwiz
8 Results on MME
Our evaluation on MME [fu2024mmecomprehensiveevaluationbenchmark] dataset is presented in Tab. 6. Our method achieves better overall (Total) results and equal or improved performance in existence and counting, demonstrating its effectiveness in object recognition. However, it does not improve position accuracy and exhibits varying behavior on color. This diversity may stem from the inherent capabilities of MLLMs, which cannot be solely enhanced through token influence balancing.
@@ 8 Results on MME
9 Other Results of POPE
Table 7: More Results on POPE [li2023evaluating].
@@ 9 Other Results of POPE
10 Different Sampling Strategies
Tab. 8 presents an ablation study on sampling strategies (non-greedy vs. greedy). We follow the non-greedy sampling setting of VCD [leng2024mitigating], where both top-p and temperature are set to 1. As shown, our method consistently improves performance across both sampling strategies.
@@ 10 Different Sampling Strategies
11 Different Model Size
@@
12 Revisiting the Accuracy–Informativeness Trade-off
MLLM Method $C_{s}$ $\downarrow$ $C_{i}$ $\downarrow$ R $\uparrow$ Len $\uparrow$
LLaVA-v1.5 base 48.8 13.4 78.6 99.8
PAI [liu2024paying] 24.8 6.9 - -
Middle [jiang2025devils] 25.0 6.7 - -
Ours_ES_12% 23.5 6.5 55.0 54.1
13 Grouping Multiple Influential Tokens with Respect to the Anchor Object
We also explored top-$2$ variants as a straightforward extension, but they produced more diffuse attributions and did not yield consistent improvements over the top-$1$ design.
Table 11: Top-k Anchor-Object Influential Token Strategies on the POPE MSCOCO Adversarial Dataset
@@ 13 Grouping Multiple Influential Tokens with Respect to the Anchor Object
top-11 83.5 82.1 82.5 82.1 84.2 83.7
top-22 83.1 81.9 79.6 80.6 81.3 80.9
14 Hyper parameter Study
Maximum $\alpha_{m}$. To determine the optimal maximum amplification factor ($\alpha_{m}$) and understand its impact on the model performance, we conducted a search over values from 1 to 6 using the LLaVA-v1.5 model on the POPE dataset for discriminative tasks. For open-ended generation tasks on a subset of the MSCOCO dataset following [deng2024seeing], we observed garbled text when $\alpha_{m}$ was set to 5; therefore, we limited our search to values from 1 to 4. As shown in Tab. 12 and Tab. 13, discriminative task on POPE is less sensitive to the value of $\alpha_{m}$. The performance on the generative task gets improved while $\alpha_{m}$ increases but later drops. Therefore, in our experiments, the optimal maximum amplification factor ($\alpha_{m}$) is set to 5 and 3 for discriminative and generation tasks, respectively.
Table 12: $\alpha_{m}$ Study For Discriminative Task On POPE [li2023evaluating] in MSCOCO Adversarial Setting.
@@ 14 Hyper parameter Study
Early Stopping Threshold. To set the early stopping threshold properly, we conducted a search on a subset of the MSCOCO dataset subset following [deng2024seeing]. Recognizing that the visual influence ratio varies across models, we first analyze the sample-wise visual influence to identify an appropriate range for this study. By searching over the corresponding range, we show results in Tab. 14. These results demonstrate that varying the ES threshold primarily mediates the trade-off between recall and hallucination rate. Our goal is to have balanced recall ($R$) and instance-level hallucination ($C_{I}$), leading us to select thresholds of $7\%$ for LLaVA-v1.5 and LLaVA-v1.6, $25\%$ for InstructBLIP, $2.5\%$ for mPLUG-Owl2 and $10\%$ for InternVL2. We additionally ran an experiment to measure the ES activation rate using LLaVA-v1.5 with ES threshold $7\%$. As shown in Tab. 15, ES fires on only $8.7\%$ of the test samples, and when it does, the generated responses are on average just 0.7 tokens shorter. This indicates that ES rarely, and only minimally, truncates outputs.
@@
15 Confidence and Visual Influence
@@
16 Question Category Results on the AMBER Dataset
@@
17 Influence Ratio in VQA
@@
18 Influence Ratio in Image Caption
@@
19 Qualitative Example on Occluded Images
@@
We include a qualitative example in Fig. 4, where a sedan and a building are partially occluded by a white truck, our method prevents the baseline model from hallucinating of persons and vehicles behind the occluding object. This demonstrates our method remains effective on images consisting of occlusions. Image caption w/o Ours: The image features a white truck parked on a city street, with graffiti covering its side. The truck is positioned near a crosswalk, and there are several other vehicles in the scene, including a car and a bus. In addition to the vehicles, there are a few people walking around the area. Image caption w Ours: The image features a large white truck parked on a city street, with graffiti covering its side. The truck is positioned near a crosswalk, and there is another vehicle visible in the background. Additionally, there are buildings in the scene, suggesting an urban setting.
20 Additional Implementation and Experimental Details
Noun tokens are identified using the spaCy library via its en_core_web_sm model. For experiments on the Amber dataset [wang2023llm], we adopt the original data splits and evaluation metrics. In the MSCOCO [lin2014microsoft] subset, we follow the data partitioning and evaluation protocol of Deng et al. [deng2024seeing], with splits available in their official repository. For the LLaVA-QA90 [liu2024visual], MME [fu2024mmecomprehensiveevaluationbenchmark] and POPE [li2023evaluating] datasets, our setup replicates that of Leng et al. [leng2024mitigating] and use their provided scoring scripts for LLaVA-QA90. Experiments on MMBench [liu2025mmbench], MM-Vet [yu2023mm] follow the VLMEvalKit_InternVL2_5 repository. All comparison methods are executed using their official code; we only modify them to enforce greedy sampling and a uniform maximum generation length to align with our experimental settings.
21 Computational Cost
In this section, we analyze the computational cost of our method. The visual encoder is executed only once, and the second pass operates on a small set of tokens. On the POPE dataset, this results in a $101.44\%$ increase in average TFLOPs, which is consistent with decoding-based approaches that rely on two forward passes. For comparison, VCD incurs an overhead of $100.10\%$.
Method TFLOPs Relative Increase
LLaVA-v1.5 $9.68$
VCD $19.37$ $100.10\%$
Ours $19.49$ $101.44\%$
22 Broader Impacts